twice the sum of the number N and 7, increased by 5 is 35
Answer by JulietG(1812)

2(N+7) + 5 = 35
Distribute
2N + 14 + 5 = 35
Add
2N + 19 = 35
Subtract 19 from each side
2N = 16
Divide each side by 2
N = 8